OpenStudy (anonymous):

Multiply the following polynomials. (3x²- 5x + 2)(7x + 1)

OpenStudy (anonymous):

(3x^2 - 5x + 2)(7x + 1) = (3x^2 * 7x) + (3x^2 * 1) + (-5x * 7x) + (-5x * 1) + (2 * 7x) + (2 * 1) = 21x^3 + 3x^2 - 35x^2 - 5x + 14x + 2 = 21x^3 - 32x^2 + 9x + 2

OpenStudy (anonymous):

expand the brackets and simplify your answer should be (21x^3-32x^2+2)

OpenStudy (anonymous):

thank you sooo much

OpenStudy (anonymous):

Distribute out the terms considering all the possible combinations:\[(\color{red} a+\color{red}b+\color{red}c)(\color{blue}d+\color{blue}e)=\color{red}a\color{blue}d+\color{red}b\color{blue}d+\color{red}c\color{blue}d+\color{red}a\color{blue}e+\color{red}b\color{blue}e+\color{red}c\color{blue}e\]